# Recursive Gaussian Process State Space Model

### v3

@@ 3 Bayesian Update Equation for Online GPSSMs
This section derives a Bayesian update equation for online GPSSMs with minimal approximations to achieve recursive learning in an arbitrarily operating domain.
@@ 3.1 Two-Step Inference
The online inference of GPSSMs is essentially a filtering problem for the hidden Markov model; therefore, most methods are based on filtering techniques, such as the extended Kalman filter (EKF) and particle filter (PF). Considering that the PF will exacerbate the computational cost issues due to GP and its particle degradation problem, we choose the EKF in this paper. The widely adopted EKF is a two-step version (one-step version can be seen in [49]), which is more numerically stable [50] and can handle irregular measurements. In each update, the two-step EKF propagates the state distribution and then corrects it with the current measurement, which are known as the prediction (time update) and correction (measurement update) steps, respectively. In the context of GPSSMs, considering the coupling between the system state and the GP, the distribution to be propagated and corrected is the joint distribution of both. These two operations can be expressed by:
@@
where $p(\bm{x}_{t+1}|\bm{x}_{t},\bm{f})$ and $p(\bm{y}_{t+1}|\bm{x}_{t+1})$ are the transition and measurement densities as defined in (4), $p(\bm{x}_{t},\bm{f}|\bm{y}_{1:t})$ denotes the joint posterior at time $t$, and $p(\bm{x}_{t+1},\bm{f}|\bm{y}_{1:t})$ is the joint prior or predicted distribution at time $t+1$. To evaluate the distribution in equations (5) and (6), the key challenges are the nonparametric nature of the GP and the nonlinearity in the transition and measurement models. Existing methods [41, 40, 42, 43, 44] typically first address the former by pre-parameterizing the GP into a fixed structure and then applying nonlinear filtering to handle the latter. However, pre-parameterization limits the operating domain, thus reducing the algorithm’s flexibility. To overcome this, we reverse the process: first handling the nonlinearity without any GP approximation, then addressing the nonparametric nature. We achieve the first procedure in the next subsection, resulting in an update equation with a factorized approximate distribution.
@@ 3.2 Factorized Approximate Distribution
To address the nonlinearity within the transition and measurement models, the following theorem is derived in this work by applying first-order linearization. To the best of our knowledge, this is the first time such a result has been derived.
Theorem 1
By approximating the transition and measurement models with first-order linearization, the approximate joint distribution between the state and GP, i.e. $q(\bm{x}_{t},\bm{f})\approx p(\bm{x}_{t},\bm{f})$ in the prediction step (5) and correction step (6) follows the factorized form:
@@
where $\bm{u}\triangleq\{u_{i}\}_{i=1}^{n_{u}}$ is a collection of a finite number of inducing points, $\bm{f}_{\neq u}$ is the function values of the function $f(\cdot)$ except for $\bm{u}$, and $p(\bm{f}_{\neq u}|\bm{u})$ is the exact GP conditional distribution similar to (2), namely,
$p(\bm{f}_{\neq u}|\bm{u})=\mathcal{N}\left(\bm{K}_{f_{\neq u}u}\bm{K}_{uu}^{-1}\bm{u},\,\bm{K}_{f_{\neq u}f_{\neq u}}-\bm{K}_{f_{\neq u}u}\bm{K}_{uu}^{-1}\bm{K}_{uf_{\neq u}}\right)$ (8)
and $q(\bm{x}_{t},\bm{u})$ is a Gaussian distribution.
Given Theorem 1, since $p(\bm{f}_{\neq u}|\bm{u})$ is known, we only need to update the finite-dimensional distribution $q(\bm{x}_{t},\bm{u})$, which build a foundation for addressing the non-parametric nature of the GP. Moreover, the approximate joint distribution remains Gaussian, which can be evaluated using closed-form update equations. In the following, we will prove Theorem 1 and then derive these equations.
@@
In this expression, the moments of the joint distribution are distinguished based on the state $\bm{x}_{t}$ and inducing points $\bm{u}$ or function values $\bm{f}$. For $q_{t}(\bm{x}_{t},\bm{u})$, the vectors $\bm{\mu}_{t}=\mathbb{E}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t}]$ and $\bm{m}_{u}=\mathbb{E}_{q(\bm{x}_{t},\bm{u})}[\bm{u}]$ represent the means of the state $\bm{x}_{t}$ and the inducing points $\bm{u}$, respectively. $\bm{P}_{t}=\operatorname{Var}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t}]$ denotes the autocovariance of $\bm{x}_{t}$, while $\bm{V}_{xu}=\operatorname{Cov}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t},\bm{u}]$ represents the cross-covariance between $\bm{x}_{t}$ and $\bm{u}$, and $\bm{S}_{uu}=\operatorname{Var}_{q(\bm{x}_{t},\bm{u})}[\bm{u}]$ signifies the autocovariance of $\bm{u}$. For the moments of $q_{t}(\bm{x}_{t},\bm{f})$, its notation follow the same logic, and the values are fully specified by the moments of $q(\bm{x}_{t},\bm{u})$ given (7), specifically:
@@
where we use the shorthand notation $\bm{k}_{fu}=\bm{K}_{fu}\bm{K}_{uu}^{-1}$. This equation can be easily derived by using (7) and the conditional distribution $p(\bm{f}|\bm{u})$ (8).
By means of the above notation, we will illustrate the proof of Theorem 1, which is based on mathematical induction. First, it can be demonstrated that at time 0, the joint distribution naturally satisfies the factorized form (7):
@@
where $q(\bm{x}_{0},\bm{u})\triangleq p(\bm{u})p(\bm{x}_{0})$, $p(\bm{f})$ denotes the GP prior over all function values $\bm{f}$, and $\bm{u}$ can be any collection of function values or an empty set. Therefore, the next step is to prove that the approximate updates of the prediction step (5) and the correction step (6) can preserve the factorized form.
For the prediction step, to address the nonlinearity in the transition model (4c), we employ first-order linearization as in the EKF, specifically linearizing the transition model as follows:
@@
where we use the notation for moments in (9) and the shorthand notation $f_{t}=f(\bm{\mu}_{t})$, $m_{f_{t}}=\mathbb{E}_{q(\bm{x}_{t},\bm{f})}[f_{t}]$, $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$. Additionally, the two Jacobian matrices in the equation are
@@
Note that, we use the hat notation $q(\bm{x}_{t+1}|\bm{x}_{t},f_{t})$ in (12) to denote the linearized model. As shown in (12), this linearized transition model depends solely on the function value $f_{t}$. Therefore, if we incorporate $f_{t}$ into the inducing-point set, i.e., $\bar{\bm{u}}=[\bm{u}^{T},f_{t}^{T}]^{T}$, the prediction result in (5) by using the linearized model (12) retains the factorized form (7):
@@
where we use the superscript "-" to denote predicted distributions, and $q^{-}(\bm{x}_{t+1},\bar{\bm{u}})=\int q(\bm{x}_{t},\bar{\bm{u}})q(\bm{x}_{t+1}|\bm{x}_{t},f_{t})\mathrm{d}\bm{x}_{t}$. It can be observed that the factorized form is preserved at the cost of augmenting the inducing-point set, an inevitable consequence of the non-parametric nature of the GP. This increase in the number of inducing points will lead to a rapid rise in computational burden, rendering inference intractable, which will be addressed in Section 4.1.
For the correction step, the derivation is similar. To handle the nonlinearity in the measurement model (4d), the first-order linearization is used again:
@@
where $\bm{\mu}_{t+1}^{-}=\mathbb{E}_{q^{-}(\bm{x}_{t+1},\bar{\bm{u}})}[\bm{x}_{t+1}]$ denotes the predicted mean of state $\bm{x}_{t+1}$ and the measurement Jacobian is $\bm{C}_{x}=\partial\bm{g}(\bm{x})/\partial\bm{x}\big{|}_{\bm{\mu}_{t+1}^{-}}$. By incorporating the linearized measurement model $q(\bm{y}_{t+1}|\bm{x}_{t+1})$ into the correction step (6), the factorized form can be retained as follows:
@@
where $q(\bm{x}_{t+1},\bar{\bm{u}})\propto q^{-}(\bm{x}_{t+1},\bar{\bm{u}})q(\bm{y}_{t+1}|\bm{x}_{t+1})$.
Thus far, we have proven Theorem 1, in which only first-order linearization is used and the inducing-point set is continually expanded to match the data distribution. In the derivation, we also obtain the approximate prediction equation (14) and correction equation (16) for the joint distribution $q(\bm{x}_{t},\bm{f})$, which can be used to derive practical update equations for it. It is evident that the joint distribution $q(\bm{x}_{t},\bm{u})$ contains the essential information of $q(\bm{x}_{t},\bm{f})$ and thus is the only component that requires updating. Given that $q(\bm{x}_{t},\bm{u})$ is Gaussian, we only need to derive its moment matching equations, which are presented in the next subsection.
@@ 3.3 Moment Matching Equation
For the prediction step, given the update equation (14), we should first add the point $f_{t}$ into the inducing point set $\bm{u}$, thus obtaining the new set $\bar{\bm{u}}$. Therefore, we define the new augmented state $\bar{\bm{X}}_{t}=[\bm{x}_{t}^{T},\bar{\bm{u}}^{T}]^{T}$, similar to $\bm{X}_{t}$. Correspondingly, we have the approximate distribution $q(\bar{\bm{X}}_{t})$ with mean $\bar{\bm{\xi}}_{t}$ and covariance $\bar{\bm{\Sigma}}_{t}$, which can be evaluated using (9) and (10) by letting $\bm{f}=\bar{\bm{u}}$. Then, by applying the update equation (14), the moment matching equation for the prediction step can be easily derived, with a result similar to that of the EKF:
@@
$\bar{\bm{\Sigma}}_{t+1}^{-}=\bm{\Phi}\bar{\bm{\Sigma}}_{t}\bm{\Phi}^{T}+\bm{\Sigma}_{f,\bar{X}}$
@@
$\bm{\Sigma}_{f,\bar{X}}=\begin{bmatrix}\bm{\Sigma}_{p}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}$
where $\bar{\bm{\xi}}_{t+1}^{-}$ and $\bar{\bm{\Sigma}}_{t+1}^{-}$ represent the predicted mean and covariance for the augmented state $\bar{\bm{X}}_{t+1}$. $\bm{\Phi}$ denotes the transition Jacobian matrix for $\bar{\bm{X}}_{t+1}$, $\bm{\Sigma}_{f,\bar{X}}$ is the process noise covariance for $\bar{\bm{X}}_{t+1}$, and $\bm{I}_{n_{u}}$ is the identity matrix.
For the correction step, the moment matching equation can be easily obtained using the approximate correction equation in (16), with the result as follows:
@@
where $\bm{P}_{t+1}^{-}$ denotes the predicted variance of the system state $\bm{x}_{t+1}$, $\bar{\bm{\xi}}_{t+1}$ and $\bar{\bm{\Sigma}}_{t+1}$ represent the posterior mean and covariance for the augmented state $\bar{\bm{X}}_{t+1}$, $\bm{H}$ is its measurement Jacobian, and $\bm{G}$ corresponds to the Kalman gain. Therefore, through the moment matching equations (17) and (18), closed-form recursive inference for online GPSSMs can be achieved.
@@
In theory, the derived equations have enabled recursive learning for GPSSMs without being restricted by the operating domain. However, as discussed in Section 3.2, the increasing number of inducing points significantly raises the computational burden for moment evaluation, which will make inference intractable. In the next section, an approximation method will be introduced to address this issue.

### v4

@@ 3 Bayesian Update Equation for Online GPSSMs
This section derives a Bayesian update equation for online GPSSMs with minimal approximations to achieve recursive learning in an arbitrary operating domain.
@@ 3.1 Two-Step Inference
The online inference of GPSSMs is essentially a filtering problem for the hidden Markov model; therefore, most methods are based on filtering techniques, such as the extended Kalman filter (EKF) and particle filter (PF). Considering that the PF not only exacerbates the computational cost of the GP but also suffers from particle degeneracy, we adopt the EKF in this paper. The widely adopted EKF is a two-step version (one-step version can be seen in [35]), which is more numerically stable [36] and can handle irregular measurements. In each update, the two-step EKF propagates the state distribution and then corrects it with the current measurement, which are known as the prediction (time update) and correction (measurement update) steps, respectively. In the context of GPSSMs, considering the coupling between the system state $\bm{x}$ and the GP, the distribution to be propagated and corrected is the joint distribution of both. These two operations can be expressed by:
@@
where $p(\bm{x}_{t+1}|\bm{x}_{t},\bm{f})$ and $p(\bm{y}_{t+1}|\bm{x}_{t+1})$ are the transition and measurement densities as defined in (4). Additionally, $p(\bm{x}_{t},\bm{f}|\bm{y}_{1:t})$ denotes the joint posterior at time $t$, and $p(\bm{x}_{t+1},\bm{f}|\bm{y}_{1:t})$ is the joint prior or predicted distribution at time $t+1$. To evaluate the distribution in equations (5) and (6), the main challenges are the nonparametric nature of the GP (that is, $\bm{f}$ is infinite dimensional) and the nonlinearity in the transition and measurement models. Existing methods [8, 27, 28, 29, 30] typically first address the former by pre-parameterizing the GP into a fixed structure and then applying nonlinear filtering to handle the latter. However, the pre-parameterization restricts the operating domain of the model (e.g., learning is limited to a predefined range such as [1,1][-1,1] in the one-dimensional input domain), reducing the flexibility of the algorithm. To overcome this, we reverse the process: first handling the nonlinearity without any GP approximation, then addressing the nonparametric nature. We achieve the first procedure in the next subsection, resulting in an update equation with a factorized approximate distribution.
@@ 3.2 Factorized Approximate Distribution
To address the nonlinearity within the transition and measurement models, the following theorem is derived in this work by applying first-order linearization.
Definition 0 (First-order linearization).
Consider a Gaussian probabilistic model
$p(\bm{a}\mid\bm{b})=\mathcal{N}\big(\bm{a}\mid\bm{h}(\bm{b}),\,\bm{S}\big),$
where $\bm{h}(\cdot)$ is a differentiable mapping and $\bm{S}$ is the covariance matrix. The first-order linearization of this model around the mean $\bm{m}_{b}$ is defined as
$p(\bm{a}\mid\bm{b})\approx\mathcal{N}\!\left(\bm{a}\,\middle|\,\bm{h}(\bm{m}_{b})+\frac{\partial\bm{h}}{\partial\bm{b}}\bigg|_{\bm{m}_{b}}(\bm{b}-\bm{m}_{b}),\,\bm{S}\right).$
Theorem 1.
Let $q(\bm{x}_{t},\bm{f})$ denote an approximation to the joint distribution $p(\bm{x}_{t},\bm{f})$ obtained in the prediction step (5) and the correction step (6), where, for brevity, $p(\bm{x}_{t},\bm{f})$ represents either $p(\bm{x}_{t+1},\bm{f}\mid\bm{y}_{1:t})$ or $p(\bm{x}_{t+1},\bm{f}\mid\bm{y}_{1:t+1})$. Let $\bm{u}\triangleq\{u_{i}\}_{i=1}^{n_{u}}$ denote a finite set of inducing points, where each inducing point $u_{i}=f(\bm{x}_{i})$ is the function value at the input location $\bm{x}_{i}$. The set $\bm{u}$ can be initialized as an arbitrary finite set or as the empty set.
In each prediction step and correction step, by applying the first-order linearization described in Definition 3.2 to both the transition model (4c) and the measurement model (4d), and by augmenting the set of inducing points $\bm{u}$ with the function value $\bm{f}(\mathbb{E}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t}])$, the approximate distribution $q(\bm{x}_{t},\bm{f})$ takes the following factorized form:
@@
where $\bm{f}_{\neq u}=\bm{f}\setminus\bm{u}$ denotes all function values except $\bm{u}$, and $p(\bm{f}_{\neq u}|\bm{u})$ denotes the GP prior conditional, analogous to (2):
$p(\bm{f}_{\neq u}|\bm{u})$ (8)
$=\mathcal{N}\left(\bm{K}_{f_{\neq u}u}\bm{K}_{uu}^{-1}\bm{u},\,\bm{K}_{f_{\neq u}f_{\neq u}}-\bm{K}_{f_{\neq u}u}\bm{K}_{uu}^{-1}\bm{K}_{uf_{\neq u}}\right)$
Furthermore, $q(\bm{x}_{t},\bm{u})$ in (7) is a Gaussian distribution.
Given Theorem 1, since $p(\bm{f}_{\neq u}|\bm{u})$ is known, we only need to update the finite-dimensional distribution $q(\bm{x}_{t},\bm{u})$, which build a foundation for addressing the non-parametric nature of the GP. Moreover, the approximate joint distribution $q(\bm{x}_{t},\bm{f})$ remains Gaussian, which can be evaluated using closed-form update equations. In the following, we will prove Theorem 1 and then derive these equations.
{pf}
@@
In this expression, the moments of the joint distribution are distinguished based on the state $\bm{x}_{t}$ and inducing points $\bm{u}$ or function values $\bm{f}$. For $q(\bm{x}_{t},\bm{u})$, the vectors $\bm{\mu}_{t}=\mathbb{E}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t}]$ and $\bm{m}_{u}=\mathbb{E}_{q(\bm{x}_{t},\bm{u})}[\bm{u}]$ represent the means of the state $\bm{x}_{t}$ and the inducing points $\bm{u}$, respectively. $\bm{P}_{t}=\operatorname{Var}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t}]$ denotes the autocovariance of $\bm{x}_{t}$, while $\bm{V}_{xu}=\operatorname{Cov}_{q(\bm{x}_{t},\bm{u})}[\bm{x}_{t},\bm{u}]$ represents the cross-covariance between $\bm{x}_{t}$ and $\bm{u}$, and $\bm{S}_{uu}=\operatorname{Var}_{q(\bm{x}_{t},\bm{u})}[\bm{u}]$ signifies the autocovariance of $\bm{u}$. For the moments of $q(\bm{x}_{t},\bm{f})$, the notations follow the same logic, and the values are fully specified by the moments of $q(\bm{x}_{t},\bm{u})$ given (7), specifically:
@@
where we use the shorthand notation $\bm{k}_{fu}=\bm{K}_{fu}\bm{K}_{uu}^{-1}$. This equation can be easily derived by using (7) and the conditional distribution $p(\bm{f}_{\neq u}|\bm{u})$ (8).
With the notation introduced above, the proof of Theorem 1 is performed using mathematical induction. First, it can be demonstrated that at time 0, the joint distribution naturally satisfies the factorized form (7):
@@
where $q(\bm{x}_{0},\bm{u})\triangleq p(\bm{u})p(\bm{x}_{0})$ is Gaussian, $p(\bm{f})$ denotes the GP prior over all function values $\bm{f}$, and $\bm{u}$ can be any collection of function values or an empty set. Therefore, the next step is to prove that, assuming after the update at time $t$ the approximate joint distribution $q(\bm{x}_{t},\bm{f})$ satisfies the factorized form (7), the approximate updates in the prediction step (5) and the correction step (6) at time $t+1$ can also preserve the factorized form (7). This is conducted in the following.
For the prediction step, to overcome the nonlinearity, the transition model (4c) is linearized around the mean of the state $\bm{x}_{t}$ and the GP as follows:
@@
where we use the notation for moments in (9) and the shorthand notation $f_{t}=f(\bm{\mu}_{t})$, $m_{f_{t}}=\mathbb{E}_{q(\bm{x}_{t},\bm{f})}[f_{t}]$, $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$. Additionally, the two Jacobian matrices in the equation are:
@@
As shown in (12), the linearized transition model $q(\bm{x}_{t+1}|\bm{x}_{t},f_{t})$ depends solely on the function value $f_{t}$. Therefore, if we include $f_{t}$ in the inducing-point set, that is, $\bar{\bm{u}}=[\bm{u}^{T},f_{t}^{T}]^{T}$, and use the linearized model in (12), the prediction result in (5) keeps the factorized form:
@@
where we use the superscript ”-” to denote predicted distributions, and $q^{-}(\bm{x}_{t+1},\bar{\bm{u}})=\int q(\bm{x}_{t},\bar{\bm{u}})q(\bm{x}_{t+1}|\bm{x}_{t},f_{t})\mathrm{d}\bm{x}_{t}$ is Gaussian. It can be observed that the factorized form is preserved at the cost of augmenting the inducing-point set, an inevitable consequence of the non-parametric nature of the GP.
For the correction step, the handling is similar. Specifically, linearize the measurement model (4d) as follows:
@@
where $\bm{\mu}_{t+1}^{-}=\mathbb{E}_{q^{-}(\bm{x}_{t+1},\bar{\bm{u}})}[\bm{x}_{t+1}]$ denotes the predicted mean of state $\bm{x}_{t+1}$ and the measurement Jacobian is $\bm{C}_{x}=\partial\bm{g}(\bm{x})/\partial\bm{x}\big|_{\bm{\mu}_{t+1}^{-}}$. By incorporating the linearized measurement model $q(\bm{y}_{t+1}|\bm{x}_{t+1})$ into the correction step (6), the factorized form can be retained:
@@
where $q(\bm{x}_{t+1},\bar{\bm{u}})\propto q^{-}(\bm{x}_{t+1},\bar{\bm{u}})q(\bm{y}_{t+1}|\bm{x}_{t+1})$ is Gaussian.
Thus far, we have proved Theorem 1 using mathematical induction. In this proof, we only use first-order linearization, and the inducing-point set is continuously expanded to match the data distribution. In contrast to existing methods [8, 27, 28, 29, 30], this result does not rely on pre-parameterizing the GP and thereby can learn over an arbitrary operating domain. Furthermore, in the derivation, we also obtain the approximate prediction equation (14) and correction equation (16) for the joint distribution $q(\bm{x}_{t},\bm{f})$, which can be used to derive practical update equations for it. It is evident that the finite dimensional distribution $q(\bm{x}_{t},\bm{u})$ contains the essential information of $q(\bm{x}_{t},\bm{f})$ and thus is the only component that requires updating. Given that $q(\bm{x}_{t},\bm{u})$ is Gaussian, we only need to derive its moment matching equations, which are presented in the next subsection.
@@ 3.3 Moment Matching Equation
For the prediction step, given the update equation (14), it includes two operations: augmenting the inducing-point set and propagating the moments. Therefore, we should first add the point $f_{t}$ into the inducing-point set $\bm{u}$, resulting into the new set $\bar{\bm{u}}$. For brevity, we define the new augmented state $\bar{\bm{X}}_{t}=[\bm{x}_{t}^{T},\bar{\bm{u}}^{T}]^{T}$, similar to $\bm{X}_{t}$. Correspondingly, we have the approximate distribution $q(\bar{\bm{X}}_{t})$ with mean $\bar{\bm{\xi}}_{t}$ and covariance $\bar{\bm{\Sigma}}_{t}$, which can be evaluated using (9) and (10) by letting $\bm{f}=\bar{\bm{u}}$. Then, by applying the update equation (14), the moment matching equation for the prediction step can be easily derived:
@@
$\bar{\bm{\Sigma}}_{t+1}^{-}=\bm{\Phi}\bar{\bm{\Sigma}}_{t}\bm{\Phi}^{T}+\bm{\Sigma}_{p,\bar{X}}$
@@
$\bm{\Sigma}_{p,\bar{X}}=\begin{bmatrix}\bm{\Sigma}_{p}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}$
where $\bm{F}_{t}$ has been defined in (12), $\bar{\bm{\xi}}_{t+1}^{-}$ and $\bar{\bm{\Sigma}}_{t+1}^{-}$ represent the predicted mean and covariance for the augmented state $\bar{\bm{X}}_{t+1}$. $\bm{\Phi}$ denotes the transition Jacobian matrix for $\bar{\bm{X}}_{t+1}$, $\bm{\Sigma}_{p,\bar{X}}$ is the process noise covariance for $\bar{\bm{X}}_{t+1}$, and $\bm{I}_{n_{u}}$ is the identity matrix.
For the correction step, the moment matching equation can be directly obtained from the approximate correction equation in (16), which takes the standard EKF form:
@@
where $\bm{C}_{x}=\partial\bm{g}(\bm{x})/\partial\bm{x}\big|_{\bm{\mu}_{t+1}^{-}}$ is the measurement Jacobian with respect to the state $\bm{x}_{t+1}$. The matrix $\bm{P}_{t+1}^{-}$ indicates the predicted variance of the system state $\bm{x}_{t+1}$. The vectors $\bar{\bm{\xi}}_{t+1}$ and $\bar{\bm{\Sigma}}_{t+1}$ are the posterior mean and covariance for the augmented state $\bar{\bm{X}}_{t+1}$. The matrix $\bm{H}$ is the measurement Jacobian with respect to the augmented state $\bar{\bm{X}}_{t+1}$, and $\bm{G}$ is the Kalman gain. Therefore, by applying the moment matching equations (17) and (18), closed-form recursive inference for online GPSSMs can be achieved in theory.
@@
However, the derived method faces a computational burden problem that prevents practical implementation. Specifically, as discussed in Section 3.2, the number of inducing points will continuously increase in each prediction step. This increase will cause the dimension of the mean vector $\bm{\xi}_{t}$ and the covariance matrix $\bm{\Sigma}_{t}$ to grow, which in turn increases the computational burden and eventually makes inference intractable. In the next section, an approximation method will be introduced to address this issue.